1) Two tiny beads are 25 cm apart with no other charges or fields present. Bead A carries 10 ?C of charge and bead B carries 1 ?C. Which one of the following statements is true about the magnitudes of the electric forces on these beads?
A) The force on A is 100 times the force on B.
B) The force on B is 10 times the force on A.
C) The force on B is 100 times the force on A.
D) The force on A is 10 times the force on B.
E) The force on A is exactly equal to the force on B.

please explain for full points
Answer #1

E) The force on A is exactly equal to the force on B.

according to newton third law each force has equal and oposite reaction hence, both particles exert same magnitude of force on each other
Answer #2

E) The force on A is exactly equal to the force on B.

F = k*q1*q2/r^2

here newton's third law is valid

the first charge exerts Force F on second charge, the second charge also exerts the same force on first charge but in opposite direction.
Answer #3

E)the force on A is exactly equal to the force B

and is equal to

F=kq1q2/r^2

F=9*10^9*10*1*10^-12/0.25^2 = 1.44N